OpenStudy (anonymous):

MIXED NUMBER 1 9/5 IN SIMPLEST FORM

OpenStudy (mathmale):

As you may know, that 1 is an integer, whereas 9/5 is a fraction with denominator 5. If you are to combine the integer, 1, with the fraction, 9/5, you'll need to convert 1 into a fraction with denom. 5. Would you try that, please?

OpenStudy (anonymous):

1 9/5 = 1 + 9/5 = 5/5 + 9/5 = 14/5